OpenStudy (anonymous):

solved by the substitution method 9x+y=72 x-6y=63 What is the solution of the problem?

OpenStudy (theviper):

x=63+6y

OpenStudy (theviper):

because x-6y=63

OpenStudy (theviper):

Put this value in the first eq

OpenStudy (theviper):

& find y then find x:)

OpenStudy (anonymous):

but what value in the first eq whats is the answer?

OpenStudy (theviper):

9x+y=72.......(1) x-6y=63 or x=63+6y Putting the value of 'x' in the eq (1) we got:- 9(63+6y)+y=72 567+54y+y=72 Now solve & find y

OpenStudy (theviper):

understood???